Sr. financial analyst, Mindy Jo, has forecasted the free cash flows of Southern Sweets to grow by 12.00% next year, 9.00% in the following year and 7.00% in the third year before leveling off to a constant growth rate of 3.00% annually. The firm’s most recently reported free cash flow was $115,000.00. The firm has $8,000.00 in marketable securities, $79,000.00 in long-term debt, $15,000.00 in preferred stock, and $52,000.00 in book common equity. Southern Sweets has 40,000.00 common shares outstanding and a cost of capital of 9.00%.

1) What is the horizon (terminal) value of Southern Sweets’ free cash flows as of year 3?

2) What is the value of operations for Southern Sweets?

3) What is the total corporate (firm) value for Southern Sweets?

4) What is the intrinsic value of common equity per share for Southern Sweets?
Answer #1

1.
=115000*1.12*1.09*1.07*1.03/(9%-3%)

=2578767.05333333

2.
=115000*(1.12/1.09)+115000*(1.12/1.09)*(1.09/1.09)+115000*(1.12/1.09)*(1.09/1.09)*(1.07/1.09)+115000*(1.12/1.09)*(1.09/1.09)*(1.07/1.09)*1.03/(9%-3%)

=2343608.56269113

3.
=115000*(1.12/1.09)+115000*(1.12/1.09)*(1.09/1.09)+115000*(1.12/1.09)*(1.09/1.09)*(1.07/1.09)+115000*(1.12/1.09)*(1.09/1.09)*(1.07/1.09)*1.03/(9%-3%)+8000

=2351608.56269113

4.
=(115000*(1.12/1.09)+115000*(1.12/1.09)*(1.09/1.09)+115000*(1.12/1.09)*(1.09/1.09)*(1.07/1.09)+115000*(1.12/1.09)*(1.09/1.09)*(1.07/1.09)*1.03/(9%-3%)+8000-79000-15000)/40000

=56.4402140672783